# TopER: Topological Embeddings in Graph Representation Learning

### v2

@@ 6 Conclusion
We have introduced a novel graph embedding method, TopER, leveraging Persistent Homology from Topological Data Analysis. TopER demonstrates strong performance in graph classification tasks, rivaling SOTA models. Furthermore, it naturally generates effective 2D visualizations of graph datasets, facilitating the identification of clusters and outliers. For future research, one promising direction is to extend TopER to temporal graph learning tasks, enabling the capture of dynamic graph trajectories that reflect evolving user behaviors over time. Another avenue worth exploring is the integration of TopER embeddings into graph foundation models, where the homogenization of graph structures could enhance the learning of transferable representations across different domains.
@@ Appendix A Further Experimental Details
A.1 Hyperparameters
Our proposed MLP algorithm is constructed with a single hidden layer. The output layer’s activation function is set to log softmax, and the loss function we used is Negative Log Likelihood Loss. The learning rate is chosen between 0.01 and 0.001. Subsequently, we investigate the impact of the number of neurons in the hidden layer, considering values from the set {16, 64, 128}, the optimizer is set to be Adam, and the number of epochs is 500. To prevent large weights and overfitting, we apply L2 regularization coefficients of 1e-3, 1e-4. The activation function for the hidden layer varies between relu, gelu, and elu. Lastly, we consider the cases of adding or not a batch normalization layer to the output of the hidden layer and set dropout values to be 0.0 or 0.5.
Table 7: Employed hyperparameters for each dataset.
Dataset Neurons Dropout Batch Norm. Decay Learning rate Activation
BZR 64 0.5 True 1e-4 0.001 gelu
COX2 128 0 True 1e-4 0.01 relu
MUTAG 16 0.5 False 1e-3 0.01 gelu
PROTEINS 64 0.5 True 1e-3 0.01 elu
IMDB-B 128 0 False 1e-3 0.001 relu
IMDB-M 16 0 False 1e-3 0.01 elu
REDDIT-B 64 0.5 False 1e-3 0.01 relu
REDDIT-5K 128 0 False 1e-3 0.01 elu
A.2 OGBG-MOLHIV Results
@@
Method Added Function Filtration Valid. AUC Test AUC
TopER-1 degree-centrality sublevel 72.76$\pm$0.23 74.44$\pm$0.20
TopER-2 atomic weight sublevel 71.89$\pm$0.12 74.25$\pm$0.16
TopER-3 Ollivier Ricci sublevel 70.11$\pm$0.28 76.79$\pm$0.24
TopER-4 Forman Ricci superlevel 71.76$\pm$0.18 78.15$\pm$0.15
TopER-5 degree superlevel 71.79$\pm$0.35 79.26$\pm$0.14
TopER-6 popularity superlevel 72.27$\pm$0.29 79.88$\pm$0.24
TopER-7 closeness sublevel 71.30$\pm$0.18 80.21$\pm$0.15
@@ A.3 Clustering Performances
In Table 9, we showcase our clustering performance across eight benchmark graph classification datasets using three widely adopted clustering metrics: Silhouette, Calinski-Harabasz, and Davies-Bouldin. These metrics serve as evaluative measures for assessing the efficacy of clustering algorithms in partitioning datasets into meaningful clusters. They gauge the degree of similarity or dissimilarity within and between clusters, offering insights into the quality of clustering outcomes. For precise definitions of Silhouette, Calinski-Harabasz, and Davies-Bouldin metrics, as well as additional details on clustering measures, refer to Gagolewski et al., (2021).
@@ B.1 Refining the point set
While we have described the main steps of TopER in Section 4, due to the repetitions of the points in $\mathcal{A}=\{(x_{i},y_{i})\}\subset\mathbb{R}^{2}$, there are some choices to be made before defining the set $\mathcal{A}$ (i.e., $\mathcal{X}$ and $\mathcal{Y}$) to get the best fitting function $L:\mathcal{X}\to\mathcal{Y}$. The main reason is that the set $\{(x_{i},y_{i})\}_{i=1}^{N}$ can contain repetitions of $x$-values ($x_{i}=x_{i+1}$), repetitions of $y$-values ($y_{i}=y_{i+1}$) or repetitions of both ($(x_{i},y_{i})=(x_{i+1},y_{i+1})$) depending on the filtration function, the threshold set $\mathcal{I}$, and the graph $\mathcal{G}$.
For the filtrations induced by node filtration functions, the number of edges can not change unless the number of nodes changes, i.e., $x_{i}=x_{i+1}\Rightarrow y_{i}=y_{i+1}$. Hence, with this elimination, we still allow keeping $y$-values the same while $x$-values are increasing. This means there can be horizontal jumps in $\mathcal{A}_{u}$. In this paper, to eliminate all horizontal jumps for filtrations with node functions, we eliminate all repetitions of $y$-values from $\mathcal{A}_{u}$. AIn particular, we remove all the points with the same $\widehat{y}$-value and add a point with a mean of $x$-values. In other words, if $y_{i}=y_{i+1}=\dots=y_{i+k}=\widehat{y}$, we define $\widehat{x}=\mbox{mean}\{x_{i},x_{i+1},\dots,x_{i+k}\}$. Then, we replace (k+1) points $\{(x_{i},\widehat{y}),(x_{i+1},\widehat{y}),\dots,(x_{i+k},\widehat{y})\}$ with one point $(\widehat{x},\widehat{y})$ in $\mathcal{A}_{u}$. This process eliminates all repetitions and horizontal jumps in $\mathcal{A}$, and we define our best-fitting line on this refined set.
@@ B.2 TopER with Alternative Quantities
While we use the most general quantities for a graph—the count of vertices and edges—in our algorithm, depending on the problem, there might be other induced quantities $(x_{i},y_{i})$ for a given subgraph $\mathcal{G}_{i}$ which can give better vectors. To keep the line-fitting approach meaningful in our model, as long as the sequences $\{x_{i}\}$ and $\{y_{i}\}$ are monotone like our node-edge counts above, for a given dataset in a domain (e.g., biochemistry, finance), one can use other domain-related quantities induced by substructure $\mathcal{G}_{i}$ as $(x_{i},y_{i})$ pair to obtain a TopER vector.
@@ B.3 Linear or Higher Order Fitting
In our experiments, we observe that linear fitting captures the growth information for node-edge pair $\{(x_{i},y_{i})\}$ well (See Figure 6), and quadratic fit and linear fit stays very close to each other. However, if one decides to use other quantities as described above and loses the monotonicity of the sequences $\{x_{i}\}$ and $\{y_{i}\}$, trying higher order fits (e.g., $y=ax^{2}+bx+c$) can be more meaningful.
@@ B.4 Interpreting TopER
Our approach involves accurately modeling the evolution of a graph throughout the filtration process. One can easily identify clusters for each class and outliers in the other datasets given in Figure 1(a) and make inferences about the different clusters and outliers. Furthermore, when the pivot $a_{f}$ is positive or negative, it can be interpreted as graph density behavior in the filtration sequence (See Figure 7).

### v3

@@ 5 Conclusion
We have introduced a novel graph embedding method, TopER, leveraging Persistent Homology from Topological Data Analysis. TopER demonstrates strong performance in graph classification tasks, rivaling SOTA models. Furthermore, it naturally generates effective 2D visualizations of graph datasets, facilitating the identification of clusters and outliers. For future research, one promising direction is to extend TopER to temporal graph learning tasks, enabling the capture of dynamic graph trajectories that reflect evolving user behaviors over time. Another avenue is the integration of TopER embeddings into graph foundation models, where the homogenization of graph structures could enhance the learning of transferable representations across different domains.
Acknowledgments.
This work was partially supported by Canadian NSERC Discovery Grant RGPIN-2020-05665: Data Science on Blockchains, National Science Foundation under grants DMS-2220613, and DMS-2229417. The authors acknowledge the Texas Advanced Computing Center (TACC) at UT Austin for providing computational resources that have contributed to the research results reported within this paper. http://www.tacc.utexas.edu.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims in the abstract and introduction accurately reflect the theoretical, methodological, and experimental contributions presented in the paper.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: The limitations are discussed in a dedicated paragraph, noting areas such as dependence on the filtration function and computational scalability to very large graphs.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. Justification: Section 4 states the full set of assumptions for each theorem, and the complete proofs are provided in the Appendix.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: The main paper and appendix include detailed descriptions of experimental setups, and an anonymous code repository is provided to ensure full reproducibility.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: The code is provided via an anonymous link in the supplementary materials, along with detailed instructions for reproducing all key experiments.
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: Training/test splits, model configurations, optimizer settings, and hyperparameter values are all specified either in the main text or in the appendix.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: All reported results include standard deviations across multiple runs, with details on how error bars were computed described in the experimental section.
  29. 8.
    Experiments compute resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: The appendix provides details about the computational environment used for all experiments, including GPU type and training time.
  33. 9.
    Code of ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: The research complies fully with the NeurIPS Code of Ethics and does not involve sensitive data, human subjects, or high-risk applications.
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: The Broader Impact section discusses potential benefits of interpretability and efficiency in scientific applications, while noting ethical considerations related to graph data usage.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: The models and datasets used in the paper do not present high risk for misuse and thus do not require additional safeguards.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: All datasets and external codebases are cited properly and used under their respective open-source licenses.
  49. 13.
    New assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: The paper introduces a new method and accompanying code, which are thoroughly documented and made available via anonymous link.
  53. 14.
    Crowdsourcing and research with human subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: The paper does not involve crowdsourcing or any research involving human subjects.
  57. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: The research does not involve human participants and does not require IRB approval.
  61. 16.
    Declaration of LLM usage
  62. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  63. Answer: [N/A]
  64. Justification: No large language models were used in the methodology or experiments; any language assistance was limited to standard editing support.
Appendix
@@ Appendix A Further Experimental Details
A.1 OGBG-MOLHIV Results
Table 7: Results for OGBG-MOLHIV of each TopERi-i.
Method Added Function Valid. AUC Test AUC
@@ A.1 OGBG-MOLHIV Results
[…]
For the OGBG-MOLHIV dataset, we further evaluated the improvements of TopER with the addition of new filtration functions. Table 7 provides the performance of each TopERi-i, where $i$ represents number of filtration functions used in the model, i.e., TopER-$i$ uses $\{(a_{1},b_{1},\dots,a_{i},b_{i})\}$ as graph embedding where $(a_{i},b_{i})$ is the pivot and growth for function $f_{i}$. We used XGBoost to rank the importance of filtration functions first, and the functions are added iteratively with this ranking. We fixed maximum tree depth =3=3, learning rates =0.035=0.035, subsample ratios =0.95=0.95, the number of estimators = 1000, and the regularization parameter lambda =45=45, where the objective function is rank:pairwise, with log loss as the evaluation metric. The seed is set to be $16$.
A.2 Time Experiments for TopER vs. PH
To compare the time efficiency and performance of TopER and persistent homology (PH),
Figure 5: Scalability. TopER run time for synthetic power law graphs [42] with node degree filtration. The mean node degree is $30$, and 100 filtration steps are used.
We conducted experiments using the same filtration function, the sublevel degree filtration. For PH, we applied Betti vectorization. Our results, summarized below, show that TopER is significantly faster than PH. Although both methods use the same filtration function, a key distinction lies in their embeddings: TopER generates 2D embeddings, whereas PH produces a vector with dimensionality equal to the number of thresholds in the filtration. Despite the considerable difference in dimensionality, TopER’s performance with 2D embeddings remains comparable to that of PH.
Figure˜5 shows that TopER scales efficiently with graph size, maintaining low runtime even with 100 filtration steps and high node degree. It processes graphs with up to 100,000 nodes in just over a minute, demonstrating its suitability for large-scale applications.
Table 8: Comparison of TopER-1 (only one filtration function) and PH in terms of time and accuracy across different datasets.
TopER-1 PH
Dataset Time Accuracy Time Accuracy # Thresholds
BZR 1.14 s 82.73 ± 2.12 5.99 s 83.70 ± 3.51 4
IMDB-B 3.27 s 73.10 ± 4.18 319.95 s 71.00 ± 4.07 65
REDDIT-B 107.65 s 79.55 ± 2.20 9173.37 s 84.50 ± 2.51 501
@@ A.3 Clustering Performances
In Table˜9, we showcase our clustering performance across eight benchmark graph classification datasets using three widely adopted clustering metrics: Silhouette, Calinski-Harabasz, and Davies-Bouldin. These metrics serve as evaluative measures for assessing the efficacy of clustering algorithms in partitioning datasets into meaningful clusters. They gauge the degree of similarity or dissimilarity within and between clusters, offering insights into the quality of clustering outcomes. For precise definitions of Silhouette, Calinski-Harabasz, and Davies-Bouldin metrics, as well as additional details on clustering measures, refer to [31].
@@
A.4 Number of Thresholds
In our experiments, we utilized a large number of thresholds to capture finer-grained information, as the model is computationally efficient and the additional cost of increasing the number of thresholds is minimal. Furthermore, in Table˜10, we evaluated the model’s performance with fewer thresholds and observed that it remains robust and highly effective even in such scenarios.
Table 10: The accuracy results of TopER with different numbers of thresholds.
# Thresholds PROTEINS REDDIT-B REDDIT-5K
[…]
A.5 Combining Filtration Functions
To assess the impact of embedding dimensions, we conducted new experiments evaluating the performance of the TopER model by progressively adding each filtration function step by step. This analysis provides insights into how the inclusion of additional filtration functions influences the model’s performance. In Table˜11, the TopER-n model represents the TopER utilizing n-filtration functions (2n features).
Table 11: Performance improvements achieved by integrating filtration functions into the TopER model. Here, TopER-n denotes the TopER model with n filtration functions.
Dataset TopER-1 TopER-2 TopER-3 TopER-4
[…]
A.6 TopER filtrations runtimes and substitute
Filtration Timing Results.
In  Table˜12, we report the computation times (in seconds) for the TopER filtration functions across various datasets. We also include the timings for the Heat Kernel Signature (HKS), which can serve as an efficient substitute for the Ollivier–Ricci curvature in certain scenarios due to its faster computation while preserving relevant structural information about the graphs. The table below summarizes the observed computation times for each filtration type across multiple benchmark datasets.
Table 12: Computation times (in seconds) of different filtration functions across datasets.
Filtration BZR COX2 MUTAG PROTEINS IMDB-B IMDB-M REDDIT-B REDDIT-5K
Degree Centrality 0.86 1.67 0.40 13.82 13.32 13.03 115.09 447.72
Popularity 0.73 1.36 0.58 5.81 13.21 15.60 111.11 414.04
Closeness 2.29 4.40 0.77 15.97 5.85 6.63 399.2 1274.00
Forman Ricci 0.80 1.56 0.58 4.63 8.29 7.38 258.32 693.18
Ollivier Ricci 130.04 121.42 48.98 313.21 277.66 428.32 1291.93 6640.95
Degree 1.14 1.16 0.61 2.62 3.27 4.22 107.65 363.32
Weight 2.25 2.80 0.90 12.89
HKS 2.07 1.94 0.56 16.90 15.54 15.72 470.01 1007.50
We evaluate the performance of our TopER model against state-of-the-art (SOTA) methods on benchmark graph classification datasets, including BZR, COX2, MUTAG, PROTEINS, IMDB-B, IMDB-M, REDDIT-B, and REDDIT-5K. Table 13 reports classification accuracy (mean ± standard deviation) across multiple runs. TopER generally achieves competitive results compared to SOTA. Ablation studies show the impact of Ricci curvature and Heat Kernel Signatures (HKS) on model performance. Notably, TopER without Ricci but with HKS recovers most of the performance lost when Ricci is removed, suggesting that HKS can serve as a viable replacement for Ollivier-Ricci curvature in capturing structural information.
Table 13: Graph classification accuracy (mean ± std) for different models across benchmark datasets. Highest scores per dataset are bold blue, second-highest are underlined blue.
Model BZR COX2 MUTAG PROTEINS IMDB-B IMDB-M REDDIT-B REDDIT-5K
SOTA 89.00±5.00 85.53±1.60 92.63±2.58 77.30±0.89 75.08±0.31 52.81±0.31 91.03±0.22 56.75±0.18
TopER 90.13±4.14 82.01±4.59 90.99±6.64 74.58±3.92 73.20±3.43 50.00±4.02 92.70±2.38 56.51±2.22
TopER w/o O. R. 87.00±4.30 77.96±8.38 87.78±7.84 74.04±3.86 73.50±3.53 50.00±5.44 91.90±2.63 56.37±1.89
TopER w/o O. R.& w/ HKS 89.63±3.65 81.58±3.54 92.08±4.23 75.20±3.59 75.00±3.49 50.67±5.58 92.75±2.47 57.33±2.02
A.7 Hyperparameters
Our proposed MLP algorithm is constructed with a single hidden layer. The output layer’s activation function is set to log softmax, and the loss function we used is Negative Log Likelihood Loss. The learning rate is chosen between 0.01 and 0.001. Subsequently, we investigate the impact of the number of neurons in the hidden layer, considering values from the set {16, 64, 128}. The optimizer is set to be Adam, and the number of epochs is 500. To prevent large weights and overfitting, we apply L2 regularization coefficients of 1e-3, 1e-4. The activation function for the hidden layer varies between ReLU, GeLU, and ELU. Lastly, we consider the cases of adding or not a batch normalization layer to the output of the hidden layer and setting dropout values to be 0.0 or 0.5. In Table˜14, we provide the details for each dataset. The last column shows the number of TopER features used for each dataset after the feature selection step.
Table 14: Employed hyperparameters for each dataset.
Dataset Neurons Dropout Batch Norm. Decay Learning rate Activation TopER Dim.
BZR 64 0.5 True 1e-4 0.001 gelu 26
COX2 128 0 True 1e-4 0.01 relu 26
MUTAG 16 0.5 False 1e-3 0.01 gelu 20
PROTEINS 64 0.5 True 1e-3 0.01 elu 26
IMDB-B 128 0 False 1e-3 0.001 relu 20
IMDB-M 16 0 False 1e-3 0.01 elu 20
REDDIT-B 64 0.5 False 1e-3 0.01 relu 24
REDDIT-5K 128 0 False 1e-3 0.01 elu 14
@@ B.1 Refining the point set
While we have described the main steps of TopER in Section˜3, due to the repetitions of the points in $\mathcal{A}=\{(x_{i},y_{i})\}\subset\mathbb{R}^{2}$, there are some choices to be made before defining the set $\mathcal{A}$ (i.e., $\mathcal{X}$ and $\mathcal{Y}$) to get the best fitting function $L:\mathcal{X}\to\mathcal{Y}$. The main reason is that the set $\{(x_{i},y_{i})\}_{i=1}^{N}$ can contain repetitions of $x$-values ($x_{i}=x_{i+1}$), repetitions of $y$-values ($y_{i}=y_{i+1}$) or repetitions of both ($(x_{i},y_{i})=(x_{i+1},y_{i+1})$) depending on the filtration function, the threshold set $\mathcal{I}$, and the graph $\mathcal{G}$.
[…]
Algorithm 1 TopER: Topological Evolution Rate
Input: Graph $\mathcal{G}$, Filtration function $f:\mathcal{V}\to\mathbb{R}$, Threshold set $\mathcal{I}=\{\epsilon_{i}\}_{i=0}^{n}$
Output: TopER vector $\mathcal{T}_{f}(\mathcal{G},\mathcal{I})$
 Initialize lists $\mathcal{X}=[]$, $\mathcal{Y}=[]$
for $i=1$ to $n$ do
  $\mathcal{G}_{i}\leftarrow$ Induced subgraph of $\mathcal{G}$ where $\mathcal{V}_{i}\subseteq f^{-1}([\epsilon_{0},\epsilon_{i}])$
  $x_{i}\leftarrow|\mathcal{V}_{i}|$
  $y_{i}\leftarrow|\mathcal{E}_{i}|$
  Append $x_{i}$ to $\mathcal{X}$
  Append $y_{i}$ to $\mathcal{Y}$
end for
 Fit a line $\mathcal{L}(x)=a+bx$ to pairs $(x_{i},y_{i})$ from $\mathcal{X}$ and $\mathcal{Y}$ using least squares
 Extract coefficients $a$ and $b$
Return $(a,b)$ as the TopER vector $\mathcal{T}_{f}(\mathcal{G},\mathcal{I})$
@@ B.2 TopER with Alternative Quantities
While we use the most general quantities for a graph—the count of vertices and edges—in our algorithm, depending on the problem, there might be other induced quantities $(x_{i},y_{i})$ for a given subgraph $\mathcal{G}_{i}$ which can give better vectors. To keep the line-fitting approach meaningful in our model, as long as the sequences $\{x_{i}\}$ and $\{y_{i}\}$ are monotone like our node-edge counts above, for a given dataset in a domain (e.g., biochemistry, finance), one can use other domain-related quantities induced by substructure $\mathcal{G}_{i}$ as a $(x_{i},y_{i})$ pair to obtain a TopER vector.
@@ B.3 Linear or Higher Order Fitting
In our experiments, we observe that linear fitting captures the growth information for node-edge pair $\{(x_{i},y_{i})\}$ well (See Figure˜6), and quadratic fit and linear fit stay very close to each other. However, if one decides to use other quantities as described above and loses the monotonicity of the sequences $\{x_{i}\}$ and $\{y_{i}\}$, trying higher order fits (e.g., $y=ax^{2}+bx+c$) can be more meaningful. In Table˜15, we present the average of the coefficients of quadratic terms when we use quadratic fit for the datasets, i.e., if we fit $y=a+bx+cx^{2}$ polynomial, we observe that the quadratic term $cx^{2}$ is mostly negligible, and tends to be a linear fit.
Table 15: Average of $x^{2}$ coefficient across datasets for quadratic fitting.
Dataset BZR COX2 MUTAG REDDIT-5k
Average of $x^{2}$ Coefficient $4.71\times 10^{-5}$ $6.61\times 10^{-4}$ $1.16\times 10^{-2}$ $1.78\times 10^{-5}$
@@ B.4 Interpreting TopER
Our approach involves accurately modeling the evolution of a graph throughout the filtration process. One can easily identify clusters for each class and outliers in the other datasets given in Figure˜1(a) and make inferences about the different clusters and outliers. Furthermore, when the pivot $a_{f}$ is positive or negative, it can be interpreted as graph density behavior in the filtration sequence (See Figure˜7).